Session: Oct 2007 (session 3)
Wavelength: 3.6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Roda at Sun Oct 28 18:10:38 2007) Success :

Onsala (Michael Lindqvist at Mon Oct 29 06:27:36 2007) Success :
RFI. A few bad TSYS/SCAN_CHECKS. Rain last hours. Otherwise no known problems.

Onsala (Michael Lindqvist at Tue Oct 30 13:08:02 2007) Minor failures :
Addition: We just found out that we have had problems with the S/X-receiver cooling. Thus, higher Tsys than normal.

Noto (P. Cassaro at Wed Oct 31 10:49:50 2007) Minor failures : RECO
Some problems during a bank change. Observation lost from 18:40 to 19:55 UT and from 01:00 to 01:45 UT.

Robledo (Cristina Garcia Miro at Wed Oct 31 19:19:28 2007) Minor failures : TSYS FS
Tsys just provided for X-band. Some RFI. Several scan_check failed. Field System (FS-9.9.0) crashed at 300/14:38UT until 300/17:32UT (14 sources lost). Received message: “shm_att: attaching memory segment failed: cannot allocate memory”. VC#14 USB with no signal because its frequency lies outside the S-Band HEMT range.

Effelsberg (A. Polatidis at Mon Nov 5 11:03:37 2007) Success :